OpenStudy (anonymous):

How do you solve systems of equations by graphing? Here are my 2 problems: 5x+2y=0 & x+2y=8

OpenStudy (anonymous):

Do you know how to find x and y intercepts?

OpenStudy (anonymous):

No, I don't think so.

OpenStudy (anonymous):

Well x intercept: when y =0 y intercept: when x=0 For the first one theyre both gonna come out intersecting the origin, and you cant make a line from that. So instead for that one lets put it in the form y=mx+c and use the slope/y intercept to graph

OpenStudy (anonymous):

Ok, but would this solve my equations? Because I have to plot each equation and then figure out where they intercept.

OpenStudy (anonymous):

Yes this will let you graph it. So for 5x+2y=0 isolate the y on the left and we get 2y=-5x+0 y=-5/2x+0 so you can see the slope is -5/2 and the y intercept is 0 Can you graph that now?

OpenStudy (anonymous):

Yes, thanks. So, on x+2y=8 , the y intercept would be 8, right? But how to you get the slope on this one?

OpenStudy (anonymous):

Would x=1?

OpenStudy (anonymous):

If you want to do that one like the last one, just get y all by its lonesome on the left. x+2y=8 2y=-x+8 y=-1/2x+4 Can you now see what the slope and y intercept is?

OpenStudy (anonymous):

The slope is -1/2 and the y intercept is 4. But how to you get the slope? I just don't get it.

OpenStudy (anonymous):

Whatever the number in front of the x is after you isolate the y on the left side is your slope x+2y=8 we subtract x from both sides 2y=-x+8 Then divide both sides by 2 y=-1/2x+4

OpenStudy (anonymous):

Oh, thanks! When you graph it, you would go up 4, then down one and over two?

OpenStudy (anonymous):

I don't follow...The y intercept on this one is 4, right? so plot (0,4) and from that point go down 1 then right 2 since your slope is -1/2. Plot that point and then draw a line through the two points

OpenStudy (anonymous):

Yeah, that is what I was saying. So, on the other one you would plot (0,0) and then go down 5 and right 3, correct?

OpenStudy (anonymous):

Down 5, right 2. And of course the point where they intersect is your solution. I think you already knew that

OpenStudy (anonymous):

Wow, thanks soooo much. I was really lost, but thanks! :D

OpenStudy (anonymous):

You're welcome

OpenStudy (anonymous):

One more question though. With 6x-10=y , how would you do that?

OpenStudy (anonymous):

How would you graph it, is that what you're asking?

OpenStudy (anonymous):

I mean how to solve. The y is already isolated, but how do you solve the rest?

OpenStudy (anonymous):

Im not understanding what you mean....do you wanna find what the slope and y intercept is?

OpenStudy (anonymous):

Alright, its just ordered different. remember the number in front of the x is the slope, and the constant is the y intercept. What do you think it would be?

OpenStudy (anonymous):

Okay... The slope would be -1/6, I think, and the y intercept is 10?

OpenStudy (anonymous):

Close...Number in front of the x is 6, so slope=6 the constant is -10 so y intercept is (0,-10)

OpenStudy (anonymous):

Ok, I'm getting the hang of it. So you would plot (0,-10) but with 6, how would you plot that?

OpenStudy (anonymous):

Think of it as 6/1 so from (0,-10) go up 6, right 1

OpenStudy (anonymous):

ok, thanks!

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

How would you solve 0.5x-4=y and -2.5x-2.5=y ?